OpenStudy (chosenmatt):

what is the difference?-2/15 -(-9/15)

OpenStudy (anonymous):

7/15

OpenStudy (anonymous):

Medals?

OpenStudy (chosenmatt):

YES

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I need a Human calculator thing

OpenStudy (anonymous):

So Ill help

OpenStudy (anonymous):

MEDALS

OpenStudy (anonymous):

NOW

OpenStudy (catlover5925):

its 7/15

OpenStudy (chosenmatt):

explaintion?

OpenStudy (catlover5925):

\[\frac{ -2 }{ 15 } +\frac{ 9 }{ 15 } = \frac{ 7 }{ 15 }\]

OpenStudy (catlover5925):

its addition because there are 2 negitives side by side

OpenStudy (chosenmatt):

@phi @ParthKohli

OpenStudy (anonymous):

-2/15 + 9/15 = 7/15 This is the correct equation just as Catlover5925 said.

OpenStudy (chosenmatt):

okay thanks

OpenStudy (anonymous):

If there is a negative next to another negative, it becomes a positive...for example, -(-9) becomes +9

OpenStudy (phi):

the first part is to know that two minuses make a plus. so change - - 9/15 to + 9/15 the second thing you need to know is that if two fractions have the *same* denominator (bottom number) then you can add (or subtract) the numerators (top number) this second rule is very important to learn.